OpenStudy (anonymous):

solve each system of equations. x+y-z=-1 x+y+z=3 3x-2y-z=-4

OpenStudy (anonymous):

any one please for the love of god explain how to do these!

OpenStudy (anonymous):

Also x+y=5 3x+z=2 4y-z=8

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Hero @jim_thompson5910

OpenStudy (anonymous):

@Psymon

OpenStudy (anonymous):

im not asking for the answers straight out. i just want to know how to do these?

OpenStudy (anonymous):

@dmezzullo

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@DDCamp

OpenStudy (ddcamp):

See how the first 2 equations look really similar, and have the same x and y coefficients (1)? Subtract equation 2 from equation 1 to get: -2z = -4 → z = 2

OpenStudy (anonymous):

yes but the z's are a negative and a positive so then what?

OpenStudy (ddcamp):

What happens when you subtract z from - z?

OpenStudy (anonymous):

no more z

OpenStudy (anonymous):

so would that make z 0?

OpenStudy (texaschic101):

x + y - z = -1 x + y + z = 3 -------------add 2x + 2y = 2 x + y + z = 3 3x - 2y - z = -4 --------------add 4x - y = -1 2x + 2y = 2 -->(-2) 4x - y = -1 ---------------- -4x - 4y = -4 (result of multiplying by -2) 4x - y = -1 ---------------add -5y = -5 y = 1 4x - y = -1 4x - 1 = -1 4x = -1 + 1 4x = 0 x = 0 x + y + z = 3 0 + 1 + z = 3 1 +z = 3 z = 3 - 1 z = 2 check... x + y - z = -1 0 + 1 - 2 = -1 -1 = -1 (correct) Any questions, feel free to ask

OpenStudy (anonymous):

Thank you so much!

OpenStudy (texaschic101):

no problem :)